OpenStudy (anonymous):

A 50-foot flagpole is at the entrance of a building that is 300 feet tall. If the length of the flagpole's shadow is 30 feet at a certain time of day, how long is the building's shadow at that time? 75 feet 150 feet 180 feet 200 feet

OpenStudy (anonymous):

@vbell134

OpenStudy (dumbcow):

set up a proportion \[\frac{x}{300} = \frac{30}{50}\] \[x = \frac{30}{50}*300 = 180\]

OpenStudy (anonymous):

180 feet

OpenStudy (anonymous):

correct thanks